OpenStudy (anonymous):

For f(x)=(1-x)/(1+x) and g(x) = (x)/(1-x) find the simplified form for f [g(x)] and state the domain. *this is an essay question so im going to need to understand every step of the problem and explain how i got the solution*** :) @ganeshie8

ganeshie8 (ganeshie8):

show what you have so far

OpenStudy (anonymous):

f(g(x)) means you plug in g(x) where ever there is an x in f(x) basically

OpenStudy (anonymous):

As this is a essay question, you should show all your steps as ganeshie required you to and we'll just point out if you're on the right path or not, because if we do this problem for you, then you'll learn nothing.

OpenStudy (anonymous):

okay so I guess i can say that I know: I take g(x) and plug it into f(X) correct?

OpenStudy (anonymous):

yes that's what you just said, so , is that what they mean by the simplified form? also what does it mean by saying: state the domain? :)

OpenStudy (anonymous):

also im not sure what the next step is after plugging in g(x) into where i see x in f(X). This is what i got: f(x) = (1- (x/1-x)) / (1+(x/1-x))

OpenStudy (anonymous):

The domain is the set of values for which the function is defined, so the function has to give you real numbers. For your case you'll be looking at where it divides by 0, and taking the square root of negative numbers that's where it will be restricted

OpenStudy (anonymous):

so okay im a little shaky on that portion, but what comes after this: f(x) = 1 - (x/1-x) / 1+(x/1-x)

ganeshie8 (ganeshie8):

that simplifies further..

OpenStudy (anonymous):

yes how do i do that

OpenStudy (anonymous):

? :)

OpenStudy (anonymous):

Well lets see what you have, I need to put it in latex haha, \[\huge f(g(x)) = \frac{ (1-\frac{ x }{ (1-x) } )}{ 1+\frac{ x }{ (1-x) } }\]

OpenStudy (anonymous):

ok yes what happens next ? :)

OpenStudy (anonymous):

i have to make the bottoms the same to subtract, so how do i do that step? :)

OpenStudy (anonymous):

There's a few ways to do this I guess, you can find a common denominator, that maybe easy but tedious >.<

OpenStudy (anonymous):

or the one in which u recommend

OpenStudy (anonymous):

Prob tedious anyway, and idk how good you are with algebra, lets just find common denominator for the numerator and denominator if that made any sense at all

OpenStudy (anonymous):

\[1-\frac{ x }{ (1-x) }\] so do this first and then for the denominator as well

OpenStudy (anonymous):

okay so do i multiply both sides with (1-x)

OpenStudy (anonymous):

\[\frac{ a }{ b } \pm \frac{ c }{ d } \implies \frac{ ad \pm bc }{ bd }\]

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Yeah, keep working on it, I think you're on the right track ;)

OpenStudy (anonymous):

okay :) f[g(x)]= f(x/1-x) = 1 - (x/1-x) / 1+(x/1-x) = [(1-x-x)/(1-x)] / [(1-x+x)/(1-x)]

ganeshie8 (ganeshie8):

simplifies further..

OpenStudy (anonymous):

^ ahahhh

OpenStudy (anonymous):

Keep working at it math

OpenStudy (anonymous):

ok :) [(1-2x/(1-x)]/[(1/1-x)]

OpenStudy (anonymous):

Then i'll divide so i'll get: -2x+1/1-x * 1-x/1 and the 1-x will cross out so i can get -2x+1 as a final answer

OpenStudy (anonymous):

That worked out pretty well

OpenStudy (anonymous):

is it correct ? :D now i have to find the domainnnn :P

OpenStudy (anonymous):

Looking at the original function, what can we say would make it undefined

OpenStudy (anonymous):

the 0 in the denominator or taking the sqrt of - numbers

OpenStudy (anonymous):

so it is x is greater than or equal to 0 because you cannot have a negative number inside a sqrt?

OpenStudy (anonymous):

actually no we dont have a sqrt lol

OpenStudy (anonymous):

so how does it divide by 0?

OpenStudy (anonymous):

No we don't so I mean like if we have this for example

OpenStudy (anonymous):

\[f(x) = \frac{ x }{ x^2-4 }\] what values of x would make this undefined

OpenStudy (anonymous):

so its x is greater than 0 :)?

OpenStudy (anonymous):

2 & -2

OpenStudy (anonymous):

\[x^2-4 \neq 0\]

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

So yours would be x cannot = 1 right?

OpenStudy (anonymous):

OOHH x cannot equal -1!!

OpenStudy (anonymous):

yess -1&1 because for one equation there is a minus sign & for another there is a plus.

ganeshie8 (ganeshie8):

-1 is not a bad value for the domain

ganeshie8 (ganeshie8):

Domain of f(g(x)) should be \[x\ne 1\]

ganeshie8 (ganeshie8):

In words : all real values except 1

ganeshie8 (ganeshie8):

you just need to look at the domain of inner function g(x) and the final simplified form

OpenStudy (anonymous):

well if i look @ the final simpliifed form, I wouldn't have a problem with the domain because my denominator is 1. correct?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ikram002p i have a question, can u explain to me why the domain here is : all reals except 1?

OpenStudy (anonymous):

@Destinymasha @Destinymasha

OpenStudy (ikram002p):

domain of g(x) sice if its 1 u would have 1/0

OpenStudy (anonymous):

why do you look @ the domain of g(x) and not the domain of f(x)

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

domain of f(x) is every where except x=-1 not 1 :)

ganeshie8 (ganeshie8):

f(g(x)) to evaluate this all it matters are g(x) and f(g(x)) you're not touching f(x) directly..